A sphere is placed inside the smallest cylinder that will contain it. If the volume of the sphere is 1.5 liters what is the volume of the space that is inside the cylinder but outside the sphere?
Answer by Boreal(15235)

volume of sphere will determine radius of cylinder as well as the height.
4/3 pi*r^3=1500 cm^3
1500*3/4/pi=r^3=7.1 cm
A cylinder with height 14.2 cm and diameter of 14.2 cm would contain something circular with radius 7.1 cm
V cylinder = pi*r^2h=pi*7.1^2*14.2=2248.82 cm^2
subtract the volume of the sphere, which is 1500 cm^3 to get 748.82 cm^2 that is inside the cylinder but outside the sphere.